# Point-DeepONet: A Deep Operator Network Integrating PointNet for Nonlinear Analysis of Non-Parametric 3D Geometries and Load Conditions

### v1

@@ Abstract
Nonlinear structural analyses in engineering often require extensive finite element simulations, limiting their applicability in design optimization, uncertainty quantification, and real-time control. Conventional deep learning surrogates, such as convolutional neural networks (CNNs), physics-informed neural networks (PINNs), and fourier neural operators (FNOs), face challenges with complex non-parametric three-dimensional (3D) geometries, directionally varying loads, and high-fidelity predictions on unstructured meshes. This work presents Point-DeepONet, an operator-learning-based surrogate that integrates PointNet into the DeepONet framework. By directly processing non-parametric point clouds and incorporating signed distance functions (SDF) for geometric context, Point-DeepONet accurately predicts three-dimensional displacement and von Mises stress fields without mesh parameterization or retraining. Trained using only about 5,000 nodes (2.5% of the original 200,000-node mesh), Point-DeepONet can still predict the entire mesh at high fidelity, achieving a coefficient of determination (R²) reaching 0.987 for displacement and 0.923 for von Mises stress under a horizontal load case. Compared to nonlinear finite element analyses that require about 19.32 minutes per case, Point-DeepONet provides predictions in mere seconds—approximately 400 times faster—while maintaining excellent scalability and accuracy with increasing dataset sizes. These findings highlight the potential of Point-DeepONet to enable rapid, high-fidelity structural analyses, ultimately supporting more effective design exploration and informed decision-making in complex engineering workflows.
@@
journal: Computer Methods in Applied Mechanics and Engineering
@@ 2.2 Data Preprocessing
Figure 4: Dataset characteristics for training and validation sets. (a) Node count distribution with violin plot insets showing spread. (b) Mass distribution with violin plot insets showing variability. (c) Sampled images from the training set by mass. (d) Sampled images from the validation set by mass.
@@ 2.3 Model Architecture
DeepONet has demonstrated its efficacy as an operator-learning framework for capturing parametric variations in complex systems, but it is constrained by using pointwise coordinates $(x,y,z)$ in the trunk network, making it challenging to encode rich global geometric information. In contrast, PointNet excels at extracting both local and global features directly from unstructured point clouds, allowing it to handle non-parametric 3D domains naturally. However, PointNet alone is not designed to learn operator mappings that relate variable load conditions to solution fields in finite element analysis (FEA) contexts. To address these complementary limitations—DeepONet’s limited global geometric embedding and PointNet’s difficulty in modeling operator mappings with changing load conditions—we introduce Point-DeepONet, a hybrid architecture that integrates the advantages of both methods.
As illustrated in Figure 7, Point-DeepONet enriches the DeepONet framework with PointNet and SIREN layers, enabling the model to handle complex 3D geometries and nonlinear responses more effectively. The branch network consists of two sub-networks: one processes the global conditions $C$ (e.g., load magnitude, direction, mass) into a latent representation $B^{(C)}\in\mathbb{R}^{B,128}$, where $B$ denotes the batch size, and the other applies PointNet operations to the point cloud $P\in\mathbb{R}^{B,N,3}$, with $N$ denoting the number of points in the finite element mesh. PointNet’s layers typically use shared multilayer perceptrons and Batch Normalization (BN) to learn robust pointwise features, and then aggregate them (e.g., via max pooling) to produce global shape descriptors. By summing the condition-based embedding $B^{(C)}$ and the point cloud-based embedding $B^{(P)}$, we obtain:
@@
On the trunk side, the input is enriched by incorporating signed distance function (SDF) values in addition to the coordinates $X=(x,y,z)$. The SDF, zero on the domain boundary and negative within the domain interior, encodes geometric context [34]. We feed $(X,SDF)\in\mathbb{R}^{B,N,4}$ into sinusoidal representation network (SIREN) layers [38], which utilize sine-based activation functions. These SIREN layers are adept at representing high-frequency variations, facilitating the capture of complex geometric details:
@@
Within the SIREN layers, the sine activation is used instead of standard activation functions, allowing effective parameterization of intricate geometric shapes. Meanwhile, in other parts of the network, we employ SiLU (Sigmoid Linear Unit) activations, defined as $\text{SiLU}(x)=x\cdot\sigma(x)$, where $\sigma$ denotes the sigmoid function. SiLU provides smooth, non-saturating nonlinearities and often leads to improved training stability and performance compared to ReLU. At the output layer, the Tanh activation function is used to ensure that predicted fields remain bounded, providing smooth outputs that range between $-1$ and $1$.
The original DeepONet architecture fuses condition-based and geometry-based information only at the final stage via a single dot product, limiting interaction between these features. Following recent advancements [39, 28, 29], we incorporate an element-wise multiplication between $B^{\alpha}$ and $T^{\alpha}$ before the final layers:
$F_{b,i,h}=B^{\alpha}_{b,h}\cdot T^{\alpha}_{b,i,h}.$ (4)
This earlier, more granular fusion encourages richer interplay between global conditions and geometric embeddings, enhancing the operator mapping’s expressiveness.
@@ 3 Results and Discussion
In this section, we present a comprehensive evaluation of three neural network architectures—PointNet, DeepONet, and our proposed Point-DeepONet—for predicting the displacement and stress fields of jet engine brackets under various load conditions. The deep learning models were trained using an NVIDIA A100 GPU with 80GB of HBM2e memory. Our objective is to assess the efficacy of each model in capturing the complex nonlinear elastoplastic behavior inherent in these structures. We begin by detailing the optimization strategies and training procedures employed to ensure a fair and consistent comparison among the models. Subsequently, we examine the influence of different resampling sizes on the accuracy of the predicted fields, highlighting how the density of sampled points affects the models’ ability to reproduce fine-grained features of the deformation and stress distributions.
@@ 3.1 Prediction Results
Model Component $u_{x}$ (mm) $u_{y}$ (mm) $u_{z}$ (mm) von Mises stress (MPa) Sampled data Full data Sampled data Full data Sampled data Full data Sampled data Full data MAE $R^{2}$ MAE $R^{2}$ MAE $R^{2}$ MAE $R^{2}$ MAE $R^{2}$ MAE $R^{2}$ MAE $R^{2}$ MAE $R^{2}$ PointNet Vertical 0.008 0.953 - - 0.003 0.927 - - 0.013 0.952 - - 11.666 0.866 - - Horizontal 0.006 0.983 - - 0.002 0.878 - - 0.008 0.980 - - 9.073 0.904 - - Diagonal 0.004 0.884 - - 0.002 0.963 - - 0.006 0.975 - - 7.639 0.859 - - DeepONet Vertical 0.026 0.626 0.018 0.731 0.007 0.709 0.006 0.800 0.048 0.654 0.031 0.834 21.842 0.592 17.289 0.691 Horizontal 0.024 0.694 0.017 0.793 0.004 0.503 0.003 0.575 0.026 0.699 0.019 0.764 20.091 0.572 16.370 0.657 Diagonal 0.006 0.578 0.006 0.542 0.005 0.753 0.004 0.781 0.016 0.776 0.013 0.807 12.555 0.645 11.172 0.706 Point-DeepONet (Our proposed) Vertical 0.007 0.964 0.006 0.979 0.003 0.923 0.003 0.949 0.012 0.961 0.011 0.981 10.541 0.882 10.139 0.881 Horizontal 0.005 0.985 0.005 0.987 0.002 0.875 0.002 0.882 0.006 0.985 0.006 0.985 7.935 0.923 7.690 0.916 Diagonal 0.003 0.903 0.003 0.890 0.002 0.953 0.002 0.955 0.005 0.976 0.005 0.978 7.090 0.873 6.866 0.876
@@
From the table, it is evident that PointNet is limited to predicting only the sampled data, as indicated by the absence of results for the full data. This constraint highlights PointNet’s scalability limitation, restricting its applicability to scenarios where the number of prediction points must match those used in training. In contrast, both DeepONet and our proposed Point-DeepONet exhibit the capability to infer over the full mesh without requiring retraining, demonstrating superior scalability and adaptability to high-resolution datasets.
@@ 3.2 Influence of Resampling Size on Displacement and Stress Distribution
To quantitatively assess the influence of resampling size, we computed the average coefficient of determination ($R^{2}$) for each model and resampling size, as shown in Figure 18(a). The results demonstrate a clear trend of increasing $R^{2}$ values with larger $N$, confirming that higher sampling densities enhance the models’ ability to explain the variance in the displacement and stress fields. Notably, the proposed Point-DeepONet generally achieved higher $R^{2}$ values compared to PointNet and DeepONet across all sampling sizes. At $N=5{,}000$, Point-DeepONet attained an average $R^{2}$ of 0.934, surpassing PointNet’s 0.925 and significantly outperforming DeepONet’s 0.758. This indicates that Point-DeepONet provides superior prediction accuracy even when using only 2.5% of the average nodes of the original geometries. Furthermore, for resampling sizes beyond $N=5{,}000$, Point-DeepONet maintained an average $R^{2}$ of 0.93 or higher, demonstrating its effectiveness in capturing the structural responses with relatively low computational overhead.
However, this improvement in accuracy comes at the cost of increased computational resources. Figure 18(b) illustrates the corresponding training times for each model and resampling size. Despite the models having a similar number of parameters, PointNet consistently required the longest training times. At $N=5{,}000$, Point-DeepONet’s training time was approximately 57 minutes, which is about 13 minutes longer than DeepONet’s 44.4 minutes but about 3 hours 36 minutes shorter than PointNet’s 4.5 hours. Importantly, Point-DeepONet achieved higher accuracy than both models, suggesting that it offers a favorable balance between training time and predictive performance. The additional training time relative to DeepONet is justified by the significant gains in accuracy, while it is more efficient than PointNet despite delivering better results.
@@ 3.3 Influence of Dataset Size on Model Performance
3.4 Effect of Input Configurations on Model Performance
In this subsection, we analyze the performance of the PointNet, DeepONet, and Point-DeepONet models under various input configurations to understand how different input parameters affect their predictive capabilities. Table 5 presents a comprehensive comparison of the models across different input configurations for predicting the displacement components ($u_{x}$, $u_{y}$, $u_{z}$) and von Mises stress under various loading conditions (Vertical, Horizontal, Diagonal). The input configurations vary in the inclusion of geometric coordinates, signed distance function (SDF), mass, force magnitude, and direction vectors.
We denote the inputs using mathematical symbols as follows: spatial coordinates are represented as $X=(x,y,z)$; the signed distance function as $SDF$; mass as $m$; force magnitude as $f$; and the direction vector as $\mathbf{d}=(d_{x},d_{y},d_{z})$. Different combinations of these inputs were used to evaluate the models’ sensitivity to various physical and geometric parameters.

### v2

# Point-DeepONet: Predicting Nonlinear Fields on Non-Parametric Geometries under Variable Load Conditions

@@ Abstract
Nonlinear structural analyses in engineering often require extensive finite element simulations, limiting their applicability in design optimization and real-time control. Conventional deep learning surrogates often struggle with complex, non-parametric three-dimensional (3D) geometries and directionally varying loads. This work presents Point-DeepONet, an operator-learning-based surrogate that integrates PointNet into the DeepONet framework to learn a mapping from non-parametric geometries and variable load conditions to physical response fields. By leveraging PointNet to learn a geometric representation from raw point clouds, our model circumvents the need for manual parameterization. This geometric embedding is then synergistically fused with load conditions within the DeepONet architecture to accurately predict three-dimensional displacement and von Mises stress fields. Trained on a large-scale dataset, Point-DeepONet demonstrates high fidelity, achieving a coefficient of determination (R²) reaching 0.987 for displacement and 0.923 for von Mises stress. Furthermore, to rigorously validate its generalization capabilities, we conducted additional experiments on unseen, randomly oriented load directions, where the model maintained exceptional accuracy. Compared to nonlinear finite element analyses that require about 19.32 minutes per case, Point-DeepONet provides predictions in mere seconds—approximately 400 times faster—while maintaining excellent scalability. These findings, validated through extensive experiments and ablation studies, highlight the potential of Point-DeepONet to enable rapid, high-fidelity structural analyses for complex engineering workflows.
@@
journal: Neural Networks
[…]
Accepted for publication in Neural Networks (2026).
DOI: 10.1016/j.neunet.2026.108560
@@ 1 Introduction
Notably, while recent methods like Geom-DeepONet [9] have advanced the field by processing point clouds, they are primarily designed for parameterized geometries where design variables (e.g., length, thickness, radius) are explicitly known and fed into the branch network. This approach is powerful for design optimization within a defined parameter space but is not applicable to non-parametric geometries, such as those resulting from topology optimization, which lack simple descriptive parameters. Furthermore, its application has been limited to variations in load magnitude. The limitations of existing methods motivated the development of Point-DeepONet. Our work addresses these critical gaps. First, by integrating PointNet, it learns a comprehensive geometric representation directly from the point cloud, eliminating the need for any parameterization. Second, it accommodates variations in both load magnitude and direction, making it a more generalized framework for real-world engineering analysis. This fundamental difference in problem definition and architectural design is why a direct quantitative comparison is challenging, and it highlights the unique contribution of our work in tackling a broader class of engineering problems.
@@ 2.2 Data Preprocessing
While Point-DeepONet demonstrates a significant speedup in analysis, it is important to consider the computational overhead of preprocessing steps, specifically the generation of the signed distance function (SDF). The SDF for each geometry is computed only once as a one-time setup cost. To provide a clear perspective on its impact, Table 2 compares the average time required for the initial FEA simulation, one-time SDF generation, and a single inference pass of our trained model.
Table 2: Quantitative comparison of computational time per geometry.
Task Average Time
Nonlinear FEA Simulation ~19.32 minutes
SDF Generation (One-time Preprocessing) ~45 seconds
Point-DeepONet Inference (Full Mesh) ~3.0 seconds
As shown in Table 2, the time required for SDF generation is negligible compared to the full FEA simulation. In practical applications such as design optimization or uncertainty quantification, where a single geometry is evaluated under numerous load conditions, the one-time SDF cost is quickly amortized. The nearly 400-fold speedup of our model refers to the inference stage, which is the critical bottleneck in such iterative analyses. Therefore, the SDF preprocessing does not undermine the overall computational efficiency of the Point-DeepONet framework.
@@
Figure 4: Dataset characteristics for training and validation sets. (a) Node count distribution with violin plot insets showing spread. (b) Mass distribution with violin plot insets showing variability. (c) Sampled images from the training set by mass. (d) Sampled images from the validation set by mass. The 20 geometries shown in (c) and (d) are assigned IDs from 1 to 20 (left to right, top to bottom) for reference in the error analysis in Section 3.5.
@@ 2.3 Model Architecture
The integration in Point-DeepONet is not a simple stacking of modules but a synergistic fusion designed to create a powerful, geometry-aware operator. The architecture, illustrated in Figure 7, leverages each component for a distinct purpose. The PointNet module in the branch network acts as a learnable feature extractor for the input geometry function, which is a significant advancement over methods that rely on fixed geometric parameters. It processes the raw point cloud $P\in\mathbb{R}^{B,N,3}$ to produce a global geometric embedding $B^{(P)}\in\mathbb{R}^{B,128}$. This embedding captures the holistic shape of the object. Concurrently, a separate MLP-based sub-network processes the global physical conditions $C$ (e.g., load magnitude, direction, mass) into a condition embedding $B^{(C)}\in\mathbb{R}^{B,128}$. These two latent vectors are then combined to form a unified context vector for the branch network:
@@
On the trunk side, the input is enriched by incorporating signed distance function (SDF) values in addition to the coordinates $X=(x,y,z)$. We feed $(X,SDF)\in\mathbb{R}^{B,N,4}$ into sinusoidal representation network (SIREN) layers [33], which utilize sine-based activation functions to effectively represent high-frequency details in the solution field:
@@
The key innovation lies in the fusion mechanism. The original DeepONet architecture fuses information only at the final stage via a single dot product, limiting the interaction between global and local features. Our approach, inspired by recent advancements [37, 10], incorporates an element-wise multiplication between the branch context vector $B^{\alpha}$ and the trunk’s point-wise feature vector $T^{\alpha}$. This operation,
$F_{b,i,h}=B^{\alpha}_{b,h}\cdot T^{\alpha}_{b,i,h},$ (4)
allows the global geometric and physical context to modulate the learned basis functions at each point in space. From a physical perspective, this element-wise modulation enables the global context (geometry and load conditions encoded in $B^{\alpha}$) to locally adapt the basis functions (encoded in $T^{\alpha}$) at each spatial location. In structural mechanics, stress and displacement fields depend on both global boundary conditions and local geometric features—a thin section responds differently than a thick section under the same global load. The element-wise multiplication naturally captures this physics: each component of $B^{\alpha}$ acts as a spatially-varying coefficient that scales the corresponding local basis function. This mechanism is conceptually analogous to how material properties or geometric parameters locally modify the Green’s function in classical elasticity theory [34]. In contrast, a single dot product aggregates all information into a scalar, losing the ability to represent such position-dependent modulation. This richer interaction allows the model to better approximate the true physical operator, particularly in problems with strong local-global coupling such as stress concentration near geometric discontinuities.
@@ 3 Results and Discussion
In this section, we present a comprehensive evaluation of our proposed Point-DeepONet architecture. The models were trained using an NVIDIA A100 GPU with 80GB of HBM2e memory. Our objective is to assess the efficacy of each model in capturing the complex nonlinear elastoplastic behavior inherent in these structures. We begin by detailing the optimization strategies and training procedures employed to ensure a fair and consistent comparison among the models. We then present detailed results from our extensive validation experiments, including an ablation study to justify our architectural choices and tests on unseen load conditions to demonstrate the model’s generalization capabilities.
@@ 3.1 Prediction Results
Model Component $u_{x}$ ($m$) $u_{y}$ ($m$) $u_{z}$ (mm) von Mises stress (MPa) Sampled data Full data Sampled data Full data Sampled data Full data Sampled data Full data MAE $R^{2}$ MAE $R^{2}$ MAE $R^{2}$ MAE $R^{2}$ MAE $R^{2}$ MAE $R^{2}$ MAE $R^{2}$ MAE $R^{2}$ PointNet Vertical 0.008 0.953 - - 0.003 0.927 - - 0.013 0.952 - - 11.666 0.866 - - Horizontal 0.006 0.983 - - 0.002 0.878 - - 0.008 0.980 - - 9.073 0.904 - - Diagonal 0.004 0.884 - - 0.002 0.963 - - 0.006 0.975 - - 7.639 0.859 - - DeepONet Vertical 0.026 0.626 0.018 0.731 0.007 0.709 0.006 0.800 0.048 0.654 0.031 0.834 21.842 0.592 17.289 0.691 Horizontal 0.024 0.694 0.017 0.793 0.004 0.503 0.003 0.575 0.026 0.699 0.019 0.764 20.091 0.572 16.370 0.657 Diagonal 0.006 0.578 0.006 0.542 0.005 0.753 0.004 0.781 0.016 0.776 0.013 0.807 12.555 0.645 11.172 0.706 Point-DeepONet (Our proposed) Vertical 0.007 0.964 0.006 0.979 0.003 0.923 0.003 0.949 0.012 0.961 0.011 0.981 10.541 0.882 10.139 0.881 Horizontal 0.005 0.985 0.005 0.987 0.002 0.875 0.002 0.882 0.006 0.985 0.006 0.985 7.935 0.923 7.690 0.916 Diagonal 0.003 0.903 0.003 0.890 0.002 0.953 0.002 0.955 0.005 0.976 0.005 0.978 7.090 0.873 6.866 0.876
@@
From the table, it is evident that PointNet is limited to predicting only the sampled data, as indicated by the absence of results for the full data. This constraint highlights PointNet’s scalability limitation. In contrast, both DeepONet and our proposed Point-DeepONet exhibit the capability to infer over the full mesh without requiring retraining, demonstrating superior scalability and adaptability to high-resolution datasets.
@@
To provide a comprehensive time metric for practical deployment, we quantify the total computational cost when 500 structural evaluations are required (Table 5). This scenario is representative of iterative engineering applications including design optimization cycles, uncertainty quantification studies, load case analyses, and parametric investigations. Traditional FEA would consume approximately 161 hours (500 × 19.32 minutes) for these evaluations. In contrast, Point-DeepONet’s workflow comprises a one-time setup—dataset generation (966 hours, amortized across multiple projects), model training (0.95 hours), and per-geometry SDF generation (0.75 minutes)—followed by rapid 3.0-second inference per evaluation. The 500 evaluations thus require only 0.42 hours (25 minutes) of iterative time, yielding a 383× speedup. Importantly, the upfront costs are one-time investments reusable across different load scenarios, design variants, and uncertainty studies. For organizations conducting repeated structural analyses, the amortized cost becomes negligible, and the effective speedup approaches 400×. The break-even point occurs after approximately 50 evaluations, confirming Point-DeepONet’s cost-effectiveness for iterative engineering workflows.
Table 5: Comprehensive comparison of total computational time for 500 structural evaluations, including all setup and iterative costs.
Task Traditional FEA Point-DeepONet
Training dataset generation (3,000 cases) 966 hours 966 hours
Model training N/A 0.95 hours
SDF generation per geometry (one-time) N/A 0.75 minutes
Evaluation per iteration 19.32 minutes 3.0 seconds
Total time for 500 evaluations 161 hours 967.2 hours (one-time)
+ 0.42 hours (iterations)
Amortized time after initial setup 161 hours 0.42 hours
Speedup factor 383×
  • 1.
    Note: The one-time setup cost (dataset generation + training + SDF) is amortized across multiple analyses. For subsequent evaluations using the same trained model, only the iterative evaluation time (0.42 hours for 500 cases) is relevant.
@@
An analysis of the error plots in Figures 9-11 indicates that the largest prediction errors are consistently localized in regions of sharp geometric discontinuity and high stress concentration. This highlights a limitation in resolving high-frequency physical phenomena, a common challenge for neural operators and an important direction for future work.
@@ 3.2 Influence of Resampling Size on Displacement and Stress Distribution
To quantitatively assess the influence of resampling size, we computed the average coefficient of determination ($R^{2}$) for each model and resampling size, as shown in Figure 18(a). The results demonstrate a clear trend of increasing $R^{2}$ values with larger $N$, confirming that higher sampling densities enhance the models’ ability to explain the variance in the displacement and stress fields. Notably, the proposed Point-DeepONet generally achieved higher $R^{2}$ values compared to PointNet and DeepONet across all sampling sizes. At N=5, [^0] N=5{,}000, Point-DeepONet attained an average $R^{2}$ of 0.934, surpassing PointNet’s 0.925 and significantly outperforming DeepONet’s 0.758. This indicates that Point-DeepONet provides superior prediction accuracy even when using only 2.5% of the average nodes of the original geometries. Furthermore, for resampling sizes beyond N=5, [^0] N=5{,}000, Point-DeepONet maintained an average $R^{2}$ of 0.93 or higher, demonstrating its effectiveness in capturing the structural responses with relatively low computational overhead.
However, this improvement in accuracy comes at the cost of increased computational resources. Figure 18(b) illustrates the corresponding training times for each model and resampling size. Despite the models having a similar number of parameters, PointNet consistently required the longest training times. At N=5, [^0] N=5{,}000, Point-DeepONet’s training time was approximately 57 minutes, which is about 13 minutes longer than DeepONet’s 44.4 minutes but about 3 hours 36 minutes shorter than PointNet’s 4.5 hours. Importantly, Point-DeepONet achieved higher accuracy than both models, suggesting that it offers a favorable balance between training time and predictive performance. The additional training time relative to DeepONet is justified by the significant gains in accuracy, while it is more efficient than PointNet despite delivering better results.
@@ 3.3 Influence of Dataset Size on Model Performance
3.4 Ablation Study and Architectural Validation
To validate that our proposed architecture is a deliberately designed framework rather than a simple stacking of modules, we analyze the performance under various input and architectural configurations. This serves a dual purpose: first, to determine the optimal set of inputs, and second, to perform an ablation study that validates the contribution of key architectural components.
Table 9 presents a comprehensive performance comparison across different input configurations. The results for Point-DeepONet, when compared row by row, effectively serve as an ablation study for the input features. For instance, comparing the configuration with ‘Trunk: X, SDF‘ to the one with ‘Trunk: X‘ highlights the significant performance gain from including the signed distance function, which provides critical geometric context to the model. Overall, the configuration using all available inputs—mass, force, and direction in the branch, and coordinates with SDF in the trunk—yielded the best performance, confirming that enriching the model with both physical and detailed geometric information is crucial for accuracy.
To specifically isolate and validate the architectural contribution of the sinusoidal representation network (SIREN) layers and the fusion mechanism, we conducted additional focused experiments. First, we compared our full Point-DeepONet model against a variant where the SIREN activations in the trunk network were replaced with standard SiLU activations. The results, summarized in Table 7, demonstrate the clear advantage of using SIREN. The superior performance of the SIREN-based model underscores its effectiveness in capturing the high-frequency details inherent in complex stress and displacement fields, a task where standard activation functions often fall short.
Second, we investigated the impact of the feature fusion strategy. Our proposed model uses an element-wise multiplication to combine the branch and trunk network outputs, allowing the global context to modulate the local features. We compared this against the standard dot product fusion used in the original DeepONet. As shown in Table 8, the element-wise multiplication approach yielded a notable improvement in the average $R^{2}$ score. This suggests that the richer, point-wise interaction enabled by element-wise multiplication is more effective for capturing complex field responses than the single aggregated value produced by a dot product. These analyses, combined with the input configuration tests, confirm that the proposed Point-DeepONet architecture is not an arbitrary combination of modules but a carefully designed framework where each component plays a vital role.
Table 7: Ablation study on the effect of SIREN activations. The table shows the average $R^{2}$ score across all displacement and stress components for the best-performing input configuration.
Model Variant Average $R^{2}$ Score
Point-DeepONet (with SIREN) 0.934
Point-DeepONet (with SiLU instead of SIREN) 0.905
Table 8: Ablation study on the feature fusion mechanism. The table shows the average $R^{2}$ score for the best-performing input configuration.
Fusion Method Average $R^{2}$ Score
Element-wise Multiplication (Proposed) 0.934
Dot Product 0.912